OpenStudy (anonymous):

suppose you are playing a game of chance. If you bet $5 on a certain event, you will collect $115 (including your $5 bet) if you win. Find the odds used for determining the payoff

OpenStudy (anonymous):

even odds 1/1 would be winning $5 and getting the original $5 back so you get 10/5 and you calculate the odds from (10-5)/5 So 115 back is (115-5)/5 = 110/5 = 22/1 odds